OpenStudy (anonymous):

PLEASE HELP, I'M BEGGING, WILL MEDAL AND FAN I DON'T LIKE TO ASK QUESTIONS, BUT I HAVE TO RIGHT NOW Bill and Ron build towers using blocks. The expression 4b + 2 represents the number of blocks Bill uses. The expression 8b – 5 represents the number of blocks Ron uses. Complete the expression to represent the total number of blocks Bill and Ron use. b −

OpenStudy (anonymous):

I think it is 12b - 3 or 8b - 5. Am I right? Please help.

MsBrains (ms-brains):

Bill = 4b + 2. Ron = 8b - 5. Bill + Ron = (4b + 2) + (8b - 5). = 4b + 8b + 2 - 5. = 12b - 3. I think you are correct.

OpenStudy (anonymous):

Thank you so much! :)

MsBrains (ms-brains):

Np! Good luck. cx

OpenStudy (anonymous):

c: